OpenStudy (anonymous):

Hockey rankings. Hockey teams receive 2 points for a win and 1 point for a tie. The Wildcats once won a championship with 60 points. They won 9 more games than they tied. How many wins and how many ties did the Wildcats have?

OpenStudy (anonymous):

2w+t=60 & w-9=t

OpenStudy (anonymous):

solve the system of equations

OpenStudy (anonymous):

Let (W) equal to the number of games won Let (T) equal to the number of games tied (W=T+9) As for the points 2 times W+1timesT= 60 The expression for (W) in the 1st equation for (W) in the 2nd equation 2 times (T+9) +T=60 2T +18 +T=60 3T=42 T=14 Since (W=T+9) (W=14+9) (W=23) Checking (2timesW+1timesT=60) 2times 23+1times14=60 46+14=60 (60=60) like this

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so its right!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ty